OpenStudy (sphott51):

Cythnthia besch wants to buy a rug for a room that is 18ft wide and 26ft long. she wants to leave a uniform strip of floor around the rug. she can afford to buy 308 sq ft of carpeting. what dimensions should tthe rug have?

OpenStudy (sphott51):

Do I put it like this : (18-2x)(25-2x) = 308 ?

OpenStudy (sphott51):

or (18-x)(26-x) ?

OpenStudy (mathstudent55):

Either one will work, but correct the 25 to 26 of the first one.

OpenStudy (sphott51):

Yeah I noticed that

OpenStudy (mathstudent55):

If you use the first one, then x is the width of the strip. Then, once you know x, the dimensions of the room will by 26 - 2x and 18 - 2x. If you use the second one, keep in mind that x means twice the width of the strip. Then, once you know x, the dimensions of the room will by 26 - x and 18 - x.

OpenStudy (sphott51):

Okay so If I'll use the (18-2x)(26-2x) =308 , it'll be 4x^2 -88x +160 =0 ---> x=20, x= 2